OpenStudy (anonymous):

Let f(x): x^3 - 6x^2 + 9x When x= ..?, the graph of y=f(x) has a min. value

OpenStudy (campbell_st):

find the 1st derivative and solve for x... you will get 2 values then find the 2nd derivative test the 1st derivative solutions in the 2nd derivative to determine their nature, max. min or horizontal point of inflexion. hope it helps

OpenStudy (campbell_st):

an alternate method is to graph the curve here is a link https://www.desmos.com/calculator

OpenStudy (anonymous):

Like this? f'(x) = 0 3x² − 12x + 9 = 0 3 (x² − 4x + 3) = 0 3 (x − 1) (x − 3) = 0 x = 1, x = 3 f''(1) = 6(1) − 12 = −6 < 0 f''(3) = 6(3) − 12 = 6 > 0 But then the answer would be 3, while on the answer sheet the right answer should be 1

OpenStudy (campbell_st):

so the minimum is at x = 3 since f"(3) > 0 and max at x = 1 since f"(1) < 0 your solution is correct

OpenStudy (anonymous):

Yeah, but the correct answer should be 1. Also may I ask why we need to find the 2nd derivative too? Is it only to use the x value we got from the 1st derivative?

OpenStudy (campbell_st):

did you graph the curve? if you graph it you'll see the relative max and mins did you enter f(x) correctly.... are you sure its not f(x) = -x^3 - 6x^2 + 9x

OpenStudy (anonymous):

Ah, I'm sorry, I got the question wrong (should be the maximum value). Thank you so much for the help!!